OpenStudy (geekfromthefutur):

The function f(x) is shown below: f(x) = (2)x +1 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (2)x + 13 f(x) = (2)x + 14 f(x) = (2)x -12 f(x) = (2)x - 13

OpenStudy (geekfromthefutur):

@bakonloverk @texaschic101 plz help medal awarded :D

OpenStudy (unklerhaukus):

a straight line has the form y(x) = mx + b where b is the intercept the difference between \(y_1(x) = mx\) and \(y_2(x) = mx + b\) is that the \(y_2\) has been shifted \(\bf{up}\) by b units

OpenStudy (unklerhaukus):

oh wait is your function \[f(x) = 2^{x+1}\] [and not \( 2x+1\)?]

OpenStudy (geekfromthefutur):

yea :D

OpenStudy (anonymous):

the answer is f(x)=2x-13

OpenStudy (unklerhaukus):

If the original function is \[y = f(x) \] then a right shift of \(a\) units, is \[y= f(x-a) \] so a left shift of \(a\) units, is \[y= f(x+a) \] so shifting \[f(x) = 2^{x +1}\] by \(a\) units to the left, is \[f(x) = 2^{x +1+a}\]

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (geekfromthefutur):

ok so is tht the answer?? jw

OpenStudy (unklerhaukus):

well, your question specifies `13 units to the left` so \(a\) is 13

OpenStudy (geekfromthefutur):

oh ok so pretty much its D. f(x) = (2)x - 13

OpenStudy (unklerhaukus):

no. how did you get that?

OpenStudy (geekfromthefutur):

oh crap sorry i mean a.

OpenStudy (unklerhaukus):

shifting f(x)=2^(x+1) by a units to the left, is f(x)=2^(x+1+a) your question specifies 13 units to the left so a is 13

OpenStudy (unklerhaukus):

have you got it yet?

OpenStudy (geekfromthefutur):

yea :D it was A. right??

OpenStudy (unklerhaukus):

no

OpenStudy (geekfromthefutur):

@UnkleRhaukus ik its not A or D i just dont understand and for a wild guess i think its B

OpenStudy (unklerhaukus):

why take wild guesses? shifting a function y = f(x) by 13 units to the left, is y = f(x+13) your function is y= 2^(x+1) so shifting to the left is y = 2^((x+13)+1) = 2^(x+13+1) = 2^( ?? )

OpenStudy (geekfromthefutur):

= 2^(14)

OpenStudy (unklerhaukus):

almost but be careful x+13+1 = x + (13+1) = x + 14

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (unklerhaukus):

Do you understand?

OpenStudy (geekfromthefutur):

yea i do but i cant exactly figure out what the answer is